When people buy a fish at a pet store the cashier can check off the color of the fish as mostly red, mostly orange or mostly yellow. Currently the probability of buying a red fish is 0.31, the probability of buying an orange fish is 0.23, and the probability of buying a mostly yellow fish is 0.13 (there are colors of fish other than red, orange, and yellow). Find the probability that the purchase of a randomly selected fish is either mostly red or mostly orange.
Answer by ikleyn(52908)

.
When people buy a fish at a pet store the cashier can check off the color of the fish as mostly red,
mostly orange or mostly yellow.
Currently the probability of buying a red fish is 0.31, the probability of buying an orange fish is 0.23,
and the probability of buying a mostly yellow fish is 0.13 (there are colors of fish other than red, orange, and yellow).
Find the probability that the purchase of a randomly selected fish is either mostly red or mostly orange.
~~~~~~~~~~~~~~~~~~~~

P(either mostly red OR mostly orange) = P(mostly red) + P(mostly orange) = 0.31 + 0.23 = 0.54,


since these events are assumed to be disjoint (= mutually exclusive) in this problem.